Use the following data to answer the next two (2) questions:

Product 1

Product 2

Product 3

Direct Material Cost

$25,000

$30,000

$35,000

Direct Labor Cost

$30,000

$40,000

$50,000

Direct Labor Hours

1,200 hours

1,800 hours

2,000 hours

Factory overhead is estimated to be $40,000 and is applied based on direct labor dollars. This overhead cost is not traceable to any particular product.

  1. Factory overhead allocated to Product 2 is

  1. $13,333.

  1. $24,000.

  2. $20,000.

  3. $14,400.

  1. The total cost of Product 1 is

  1. $55,000.

  1. $62,500.

  2. $65,000.

  3. $83,333.
Answer #1

Total Labor cost = 30000+40000+50000 = $120000

a) Factory overhead allocated to product 2 = 40000/120000*40000 = $13333

So answer is a) $13333

b) Total Cost of product 1 :

Overhead applied = 40000*30000/120000 = 10000

Total product cost = 25000+30000+10000 = 65000

So answer is o) $65000